Use appropriate software to graph the surfaces
Z=7-(x^2+y^2)
and
z=1-2x+6y
on a
common screen using the domain
|x|≤10, |y|≤10
and observe the curve of intersection
of these surfaces. Show that the projection of this curve onto the xy–plane is a circle by
showing the calculation and the details of the circle equation.
1
Expert's answer
2019-04-28T17:14:24-0400

Intersection of these surfaces:


"z = 7-x^2-y^2=1-2x+6y""7+9= x^2-2x +1+y^2+6y+9""(x-1)^2+(y+3)^2=4^2"

Therefore, it is circle with radius 4, and center in point x=1, y=-3.